A solution of Na3PO4 that is 0.2000N as a salt is used to precipitate the Mg as MgNH4PO4 from a sample of dolomite containing 14.01% MgCO3. What volume is theoretically required if 2.0000g of sample is dissolved in water and diluted to 250 mL and an aliquot of 50 mL is used?


1
Expert's answer
2021-01-29T08:23:35-0500

The ionic reaction of the titration is:

Mg2+ + HPO42- + NH3 "\rightarrow" MgNH4PO4"\downarrow" .

Here, one mole of HPO42- ions react with one mole of Mg2+ ions:

"n(Mg^{+2}) = n(HPO_4^{2-})" .

It is evident that the number of the moles of HPO42- and of Na3PO4 are the same:

"n(HPO_4^{2-}) = n(Na_3PO_4)"


Therefore, the volume of Na3PO4 needed for the titration is:

"V(Na_3PO_4) = \frac{n(PO_4^{3-})}{c} = \frac{n(Mg^{2+})}{0.2000}" .


The number of the moles of Mg2+ equals to the number of the moles of MgCO3 in the aliquot:

"n(Mg^{2+}) = n(MgCO_3)" .


The number of the moles of MgCO3 is its mass divided by its molar mass 84.31 g/mol:

"n(MgCO_3) =\frac{m}{M} = \frac{m}{84.31}" .


The mass of MgCO3 in the sample of 2.0000 g is:

"m(MgCO_3, sample) = 2.0000\cdot 0.1401 = 0.2802" g.


The mass of MgCO3 in 50 mL aliquot of the solution is 250/50=5 times less than in 250 mL solution:

"m(MgCO_3, aliquot) = 0.2802\/5 = 0.05604" g.


Therefore, the number of the moles of MgCO3 in the aliquot is:

"n(MgCO_3, aliquot) = \frac{0.05604}{84.31} = 0.0006647" mol.


Finally, the theoretical volume of Na3PO4 is:

"V(Na_3PO_4) = \frac{0.0006647}{0.2000} = 0.003323" L or 3.32 mL.


Answer: 3.32 mL of Na3PO4 solution that is 0.2000N is theoretically required if 2.0000g of sampleof dolomite containing 14.01% MgCO3 is dissolved in water and diluted to 250 mL and an aliquot of 50 mL is used.